26. Dell Corp. produces three products and is currently facing a labor shortage - only 20,000 hours are available this month. The selling price, costs, labor requirements, machine requirements and demand of the three products are as follows: A. In what order should Dell prioritize (rank from 1st choice to 3rd) production of the products? To determine priority you first determine what resource is constrained (in this case it is labor hours). Then calculate Unit Contribution Margin. Using unit contribution margin, you calculate Contribution Margin per Unit of constrained resource. The product with the highest contribution margin per unit of constrained resource is the first priority. Rank the other products accordingly. ID: A Product A Product B Product C Selling price $24.00 $31.00 $50.00 Variable cost per unit $15.00 $25.00 $30.00 Direct labor hours per unit 3 1.5 2 Machine hours per unit 2 3 1 Demand 5,000 2,000 2,500 B. How many units of each product should be produced to maximize profit during the labor shortage? We will want to produce as many as are needed (demand) of the highest priority product. Next we produce as many as we need (demand) of the 2nd priority product and continue with each product until we have used all of the constrained resource. Lower priority products we will not be able to produce all of the demand.
Answer #1

Requirement 1.

Product A Product B Product C
Selling price 24 31 50
Variable cost per unit 15 25 30
Contribution per unit=(SP-VC) $9 $6 $20
Contribution/labour hour (9/3)=$3 $(6/1.5)=$4 (20/2)=$10

Hence product with maximum contribution per direct labor should be made in priority.

Hence the Rank should be C,B,A

Requirement 2.

Units of each product to be produced:

Product C = demand * direct labor hrs = 2500 * 2 = 5000 hrs So 2500 units all are produced

Product B = demand * direct labor hrs = 2000 * 1.5 = 3000 hrs So 2000 units all are produced

Product A = hours left / DL = 12000 / 3 = 4000  So 4000 units are produced